# On the emergence of numerical instabilities in Next Generation Reservoir Computing

### v1

@@ 3.4. Key example: Lorenz system with Explicit forward Euler method
(10) $\mathbf{x}_{n+1}=\mathbf{x}_{n}+h\mathbf{F}(\mathbf{x}_{n}),$
which defines a discrete-time dynamical system that approximates the continuous flow of the original dynamics for small $h$. For this time series, let us consider the NGRC model with $k=1$ (absence of any time-delayed coordinates). Figure 1 shows the attractor reconstruction and forecasting capability of the NGRC model for the Lorenz system Eq. 9. The NGRC model captures the original dynamics, successfully predicting the trajectory for up to five Lyapunov times. This also extends to other dynamical features of the original dynamics.
E. Lorenz realized that extracting successive local maxima of the $z-$component yields a one-dimensional return map that is topologically conjugated to the tent map [Lor63], which is referred to as the successive local maxima map. The left panel in Fig. 2 shows that the NGRC model successfully captures this map. The right panel of Fig. 2 displays the power spectrum density of the three variables for the original dynamics and NGRC model, showing that the NGRC model also reproduces the long-term statistics of the Lorenz attractor. These observations demonstrate NGRC’s ability to recover the chaotic dynamics using only observed trajectory data, without explicit knowledge of the underlying equations.
Figure 2. NGRC model captures topological and statistical features of the original dynamics. The left panel displays the Poincaré return map of the successive local maxima of the original dynamics (in blue) and the reconstructed one (in red). The right panel depicts the power spectrum density of the original dynamics (in blue) and the reconstructed one (in red). Both curves mostly overlap for all frequencies, indicating that the NGRC model captures correctly the long-term statistics on the attractor. The parameters are the same used for Fig. 1.
3.5. Perturbation bounds for the least squares
@@
where $\mathbf{C}=(\mathbf{c}_{1}^{\top},\dots,\mathbf{c}_{d}^{\top})$ and each row vector $\mathbf{c}_{i}^{\top}$ corresponds to the coefficient vector of the $i$-th component of the vector field in the polynomial basis, where its magnitude depends on the time step $h$  [^1] Since the explicit forward Euler method evaluates the vector field at current state only, the true coefficients $\mathbf{c}_{i}$ have zero entries corresponding to monomials functions evaluated at any of the time-delayed coordinates.. Since the problem can be solved independently for each component $i$, let us temporally drop the $i-$th dependence. If $\Psi$ is of full column rank, then the least squares solution ($\beta=0$) is unique and is attained by $\mathbf{c}$:
@@
where $\mathbf{y}$ is the vector in Eq. 4. However, in practice, numerical linear algebra is required to assess the accuracy of the computed solution. As a result, the NGRC training corresponds to solving a perturbed least squares problem of the form
(11) $\mathbf{w}=\operatorname*{arg\,min}_{\mathbf{u}\in\mathbb{R}^{m}}\big{\{}\|\mathbf{y}+\Delta\mathbf{y}-(\Psi+\Delta\Psi)\mathbf{u}\|^{2}\big{\}},$
@@
  • The magnitude of the minimum residual $\|\mathbf{y}-\Psi\mathbf{w}\|$, measuring how well the model fits the data. This introduces the closeness of fit $\theta=\sin^{-1}\big{(}{\frac{\|\mathbf{y}-\Psi\mathbf{w}\|}{\|y\|}}\big{)}$. This angle characterizes the geometric misalignment between the $\mathbf{y}$ and the range of the feature matrix [TB97].
@@
The following first-order bound quantifies numerical accuracy quantified by the relative error [GVL96, Theorem 5.3.1]: for $\varepsilon:=\max\Big{\{}\frac{\|\Delta\Psi\|}{\|\Psi\|},\frac{\|\Delta\mathbf{y}\|}{\|\mathbf{y}\|}\Big{\}}$ such that $\varepsilon\kappa(\Psi)<1$ and $\sin(\theta)\neq 1$:
(12) $\frac{\|\mathbf{w}-\mathbf{c}\|}{\|\mathbf{c}\|}\leq\varepsilon\Big{(}\frac{2\kappa(\Psi)}{\cos(\theta)}+\tan{\theta}\kappa(\Psi)^{2}\Big{)}+\mathcal{O}(\varepsilon^{2}).$ ) .
@@
Even when $\Psi$ has full column rank, a large condition number (i.e., $\Psi$ is ill-conditioned) can cause the least squares solution to deviate drastically due to the perturbation. More precisely, $\varepsilon$ above can be quantified in terms of the machine precision, which in our case is $\varepsilon_{\mathrm{machine}}=\mathcal{O}(10^{-16})$. The heuristics is that if $\varepsilon_{\mathrm{machine}}=\mathcal{O}(10^{q})$ and $\kappa(\Psi)=\mathcal{O}(10^{q^{\prime}})$, the computed solution has about $q-q^{\prime}$ correct decimal digits. Hence, we say $\Psi$ is ill-conditioned with respect to the machine precision whenever $\varepsilon_{\mathrm{machine}}\kappa(\Psi)$ is large. For instance, if the condition number grows beyond $10^{13}$, the solution can have at best digits around $0.001$. The other quantity is the closeness of fit $\theta$. If the closeness of fit is large, the relative error might scale with the second term in Eq. 12, i.e., $\kappa(\Psi)^{2}$. Consequently, for an ill-conditioned $\Psi$, the numerical accuracy can be bad.
A similar result of Eq. 12 also holds for the Tikhonov regularization ($\beta>0$), where $\frac{\sigma_{1}}{\sqrt{\beta}}$ plays the role of the condition number [Han98, 5.1.1]. By increasing $\beta$, the condition number becomes smaller, and thus the regularized solutions are less sensitive to perturbations. However, increasing $\beta$ also makes the regularization error larger.
A well-established issue in numerical analysis [GVL96, TB97, Hig02] is the numerical stability for solving the least squares problem. To compare the impact of different numerical algorithms, we select standard methods in python using scipy [VGO+20] to solve the least squares problem, including:
@@
  • Cholesky decomposition: The first method is based on solving the least squares via regularized normal equation for the unknown $\mathbf{u}$:
    (13) $(\Psi^{\top}\Psi+\beta\mathbf{1}_{m})\mathbf{u}=\Psi^{\top}\mathbf{y}_{i},\quad i=1,\dots,d,$
@@
    which is computed numerically by Cholesky factorization in scipy.linalg.solve, assuming that the matrix is positive definite. The relative error for solving linear equations is slightly different from the least squares: it scales quadratically with the condition number $\kappa(\Psi)$, because of the factor $\Psi^{\top}\Psi$, and does not depend on $\theta$ (see [GVL96, TB97, Hig02] for detailed exposition). For $\beta>0$, we expect to scale as $\sigma_{1}^{2}/\beta$.
@@
    (14) $\mathbf{w}_{i}=\Psi_{\beta}^{\dagger}\mathbf{y}_{i},$
@@
    Although this method can be computationally expensive, it is an option method to solve numerically the Tikhonov regularization [Han98]. Solving the least square problem with SVD is known to be stable [TB97, 19.4] [^2] The built-in implementation of SVD is based on a QR decomposition [ABB+99]. Solving least square problems based on QR decomposition satisfies Eq. 12. .
  • LU decomposition: The third method corresponds to solving Eq. 7 as it is mathematically written, where the inverse is computed numerically using LU decomposition via scipy.linalg.inv. This is similar to Choleskey decomposition, scaling quadractically with $\kappa(\Psi)$ for $\beta=0$ and as $\sigma_{1}^{2}/\beta$ for $\beta>0$.
Other commonly used numerical methods are pseudo-inverse and least square methods, scipy.linalg.pinv and scipy.linalg.lstsq, respectively, but those inherently add a built-in regularizer using the definition of $\varepsilon-$rank$(\Psi)$, where singular values are considered zero [GVL96, Theorem 2.5.3] whenever are upper bounded by $\sigma_{1}\times\varepsilon_{\mathrm{machine}}$.
The example of the Lorenz system generated by the explicit forward Euler method is also good in this regard. We expect that the closeness of fit is small because we would be fitting polynomial vector fields, using polynomial basis functions. As an illustration, for the parameters used in Fig. 1, the closeness of fit for these three algorithms is shown in Table 1. For this example, the condition number is $\kappa(\Psi)=8983.1469$. On the one hand, all results are of order $10^{-14}$, indicating that the fit is very close. On the other hand, there are differences among the algorithms. For this particular trajectory and parameters, SVD is the method with the minimum closeness of fit for all Lorenz’s components. As we will see in Section 6, SVD is the most robust method against the ill-conditioning of $\Psi$. For all simulations in this paper, we utilize the SVD algorithm, unless stated as in Section 6.
@@
Cholesky $3.12329\times 10^{-14}$ $8.88461\times 10^{-14}$ $1.32268\times 10^{-14}$
SVD $2.07528\times 10^{-14}$ $2.44480\times 10^{-14}$ $2.28179\times 10^{-15}$
LU $2.57541\times 10^{-14}$ $8.07677\times 10^{-14}$ $5.45502\times 10^{-14}$
@@
Table 1. Resulting closeness of fit for Lorenz system, where $\theta_{x}$, $\theta_{y}$ and $\theta_{z}$ corresponds to the quantity for each component. The parameters are the same used for Fig. 1.
4. Problem statement
As aforementioned, the accuracy of a least squares solution is strongly affected by the conditioning of the feature matrix. Our focus here is to investigate under which conditions $\Psi$ becomes ill-conditioned. Specifically, we study how the condition number varies with the delay dimension $k$, the time lag $\tau$, and the length of the training data $N_{\mathrm{train}}$. We observe that $\Psi$ possesses rich internal structure: for $k>1$, it is simultaneously Vandermonde-like and Hankel-like because each entry corresponds to a monomial evaluated along a time-delayed trajectory of the original dynamics. As we will discuss, analyzing the conditioning of $\Psi$ is non-trivial and requires combining tools from numerical linear algebra with ergodic properties of dynamical systems.
Before proceeding, it is worth highlighting why this conditioning analysis matters. While the illustrative toy example in the previous section focuses on a polynomial vector field, it sheds light on learning dynamical systems from data using NGRC.
@@ 4. Problem statement
  • Governing equations from data and NGRC stability. When training NGRC, the task reduces to finding the governing equations of a dynamical system directly from time series data [CM87, BRT94a, BRT94b]. This is an intuitive perspective into the autonomous NGRC dynamics during testing. Even small modeling errors — especially for terms that are absent in the true system — can cause instability when fitting an ODE from data, as already identified in [YB07]. For example, consider we attempt to learn $\dot{x}=-x,\dot{y}=0$ but instead recover $\dot{x}=-x,\dot{y}=\varepsilon y$. Although $\varepsilon$ may be small, the trajectory of the learned system diverges due to exponential growth in $y$. Similarly, if $\Psi$ is ill-conditioned, the estimated readout weights may differ substantially from the true ones, leading to unstable NGRC dynamics [ZdSC25].
  • Regularization is necessary for ill-conditioned $\Psi$. Our analysis also clarifies when regularization ($\beta>0$) is required. Tikhonov regularization, as used in Eq. 6, is a standard method for mitigating the effects of ill-conditioned matrices. Selecting an appropriate $\beta$ is non-trivial and has been studied extensively [Han98]. For forecasting tasks, it has been a common approach to train the model to predict forward one step, and then select the regularizer parameter $\beta$ that ensures good prediction performance over multiple time steps ahead. Our results provide a theoretical foundation for this practice by identifying regimes where regularization is not just helpful but essential.
5. The structure of $\Psi$ characterizes its conditioning

### v2

@@ II.4 Key example: Lorenz system with explicit forward Euler method
$\mathbf{x}_{n+1}=\mathbf{x}_{n}+h\mathbf{F}(\mathbf{x}_{n}),$ (10)
which defines a discrete-time dynamical system that approximates the continuous flow of the original dynamics for small $h$. Let us consider the NGRC model with $k=1$ (absence of any time-delayed coordinates). Figure 1 shows the attractor reconstruction and forecasting capability of the NGRC model for the Lorenz system Eq.˜9. The NGRC model captures the original dynamics, successfully predicting the trajectory for up to five Lyapunov times. This also extends to other dynamical features of the original dynamics.
Edward Lorenz realized that extracting successive local maxima of the $z$-coordinate yields a one-dimensional return map, which is referred to as the successive local maxima map, that is topologically conjugated to the tent map [34]. The left panel in Fig.˜2 shows that the NGRC model successfully captures this map. The right panel of Fig.˜2 displays the power spectrum density of the three variables for the original dynamics and NGRC model, showing that the NGRC model also reproduces the long-term statistics of the Lorenz attractor. These observations demonstrate NGRC’s ability to recover the chaotic dynamics using only observed trajectory data, without explicit knowledge of the underlying equations.
Figure 2: NGRC model captures topological and statistical features of the original dynamics. The left panel displays the Poincaré return map of the successive local maxima of the original dynamics (in blue) and the reconstructed one (in red). The right panel depicts the power spectrum density of the original dynamics (in blue) and the reconstructed one (in red). Both curves mostly overlap for all frequencies, indicating that the NGRC model captures correctly the long-term statistics on the attractor. The parameters are the same used for Fig.˜1.
II.5 Perturbation bounds for the least squares
@@
where $\mathbf{C}=(\mathbf{c}_{1}^{\top},\dots,\mathbf{c}_{d}^{\top})$ and each row vector $\mathbf{c}_{i}^{\top}$ corresponds to the coefficient vector of the $i$-th coordinate of the vector field in the polynomial basis, where its magnitude depends on the time step $h$  [^1] Since the explicit forward Euler method evaluates the vector field at current state only, the true coefficients $\mathbf{c}_{i}$ have zero entries corresponding to monomials functions evaluated at any of the time-delayed coordinates.. Since the problem can be solved independently for each coordinate $i$, let us temporally drop the $i$-th dependence. If $\Psi$ is of full column rank, then the least squares solution ($\beta=0$) is unique and is attained by $\mathbf{c}$:
@@
where $\mathbf{y}$ is the vector in Eq.˜4. However, in practice, numerical linear algebra is required to assess the accuracy of the computed solution. As a result, the NGRC training corresponds to solving a perturbed least squares problem of the form
$\mathbf{w}=\operatorname*{arg\,min}_{\mathbf{u}\in\mathbb{R}^{m}}\big\{\|\mathbf{y}+\Delta\mathbf{y}-(\Psi+\Delta\Psi)\mathbf{u}\|^{2}\big\},$ (11)
@@
  • The magnitude of the minimum residual $\|\mathbf{y}-\Psi\mathbf{w}\|$, measuring how well the model fits the data. This introduces the closeness of fit $\theta=\sin^{-1}\big({\frac{\|\mathbf{y}-\Psi\mathbf{w}\|}{\|y\|}}\big)$. This angle characterizes the geometric misalignment between the $\mathbf{y}$ and the range of the feature matrix [28].
@@
The following first-order bound quantifies the numerical accuracy quantified by the relative error [27, Theorem 5.3.1]: for $\varepsilon:=\max\Big\{\frac{\|\Delta\Psi\|}{\|\Psi\|},\frac{\|\Delta\mathbf{y}\|}{\|\mathbf{y}\|}\Big\}$ such that $\varepsilon\kappa(\Psi)<1$ and $\sin(\theta)\neq 1$:
[…]
@@
Even when $\Psi$ has full column rank, a large condition number (i.e., $\Psi$ is ill-conditioned) can cause the least squares solution to deviate drastically due to the perturbation. More precisely, $\varepsilon$ above can be quantified in terms of the machine precision, which in our case is $\varepsilon_{\mathrm{machine}}=\mathcal{O}(10^{-16})$. The heuristics is that if $\varepsilon_{\mathrm{machine}}=\mathcal{O}(10^{q})$ and $\kappa(\Psi)=\mathcal{O}(10^{q^{\prime}})$, the computed solution has about $q-q^{\prime}$ correct decimal digits. Hence, we say $\Psi$ is ill-conditioned with respect to the machine precision whenever $\varepsilon_{\mathrm{machine}}\kappa(\Psi)$ is large. For instance, if the condition number grows beyond $10^{13}$, the solution can have at best digits around $0.001$. The other quantity is the closeness of fit. If the closeness of fit is large, the relative error might scale with the second term in Eq.˜12, i.e., $\kappa(\Psi)^{2}$. Consequently, an ill-conditioned $\Psi$ might lead to large numerical inaccuracies. The Tikhonov regularization ($\beta>0$) is an option to avoid the ill-conditioning, where $\kappa(\Psi)$ is replaced by the regularized condition number $\kappa_{\beta}(\Psi)=\frac{\sigma_{1}}{\sqrt{\beta}}$ [36, Theorem 5.1.1]. By increasing $\beta$, the condition number becomes smaller, and thus the regularized solutions are less sensitive to perturbations.
II.6 Numerical algorithms for the regularized least squares
Different numerical algorithms are available to compute the regularized least squares solution. For ill-conditioned problems, the computed solutions can vary drastically, since the numerical stability of these algorithms depends differently on the condition number. Although this issue is well-known in numerical analysis [27, 28, 29], this has not been examined in the context of reservoir computing. To compare the impact of different solvers, we adopt three approaches used in the reservoir computing literature to solve the regularized least squares problem. All algorithms are implemented in pythonscipy [37] package.:
@@ II.6 Numerical algorithms for the regularized least squares
  • Cholesky decomposition: The first method is based on solving the least squares via the regularized normal equation for the unknown $\mathbf{u}$:
    $(\Psi^{\top}\Psi+\beta\mathbf{1}_{m})\mathbf{u}=\Psi^{\top}\mathbf{y}_{i},\quad i=1,\dots,d,$ (13)
@@
    which is computed numerically by Cholesky factorization in scipy.linalg.solve, assuming that the matrix is positive definite. The relative error for solving linear equations is slightly different from Eq.˜12: it scales quadratically with the condition number $\kappa(\Psi)$, because of the factor $\Psi^{\top}\Psi$, and does not depend on $\theta$ (see [27, 28, 29] for detailed exposition). For $\beta>0$, we expect to scale as $\sigma_{1}^{2}/\beta$.
@@
    $\mathbf{w}_{i}=\Psi_{\beta}^{\dagger}\mathbf{y}_{i},$ (14)
@@
    Although this method can be computationally expensive, it is a viable method to solve numerically the Tikhonov regularization [36]. Solving the least square problem with SVD is known to be stable [28, Theorem 19.4] [^2] The built-in implementation of SVD is based on a QR decomposition [75]. Solving least square problems based on QR decomposition satisfies Eq. 12..
  • LU decomposition: The third method corresponds to solving Eq.˜7 as it is mathematically written, where the inverse is computed numerically using LU decomposition via scipy.linalg.inv. This is similar to Choleskey decomposition, scaling quadratically with $\kappa(\Psi)$ for $\beta=0$ and as $\sigma_{1}^{2}/\beta$ for $\beta>0$.
[…]
II.7 Numerical algorithms in the ill-conditioned regime
The Lorenz system generated by the explicit Euler method provides a clear example of the interplay between hyperparameters and numerical accuracy during NGRC training. By changing a single hyperparameter, in this case the delay dimension $k$, the NGRC model can transition from a well-conditioned regime to an extremely ill-conditioned regime, with the condition number scaling as the inverse of $\varepsilon_{\mathrm{machine}}$. In this ill-conditioned regime, the solutions obtained by different numerical algorithms described in Section˜II.6 can differ by orders of magnitude.
First, consider again the case $k=1$, using the parameters from Fig.˜1. The condition number is small, $\kappa(\Psi)=149.24$, and the closeness of fit for each Lorenz coordinate across the three algorithms is shown in Table˜1. Because the NGRC model is polynomial and fits a polynomial vector field, the closeness of fit is consistently of order $10^{-14}$, indicating a high fitting accuracy. Consequently, the maximum pairwise difference between solutions as defined in
$\begin{split}\Delta=\max_{i=x,y,z}\max\Big\{\|\mathbf{w}_{i}^{\mathrm{CHO}}-\mathbf{w}_{i}^{\mathrm{SVD}}\|,\|\mathbf{w}_{i}^{\mathrm{CHO}}-\mathbf{w}_{i}^{\mathrm{LU}}\|,\\ \|\mathbf{w}_{i}^{\mathrm{SVD}}-\mathbf{w}_{i}^{\mathrm{LU}}\|\Big\},\end{split}$ (15)
is also on the order of $10^{-14}$, consistent with the error bounds predicted by Eq.˜12.
@@ II.7 Numerical algorithms in the ill-conditioned regime
Cholesky $1.51579\times 10^{-14}$ $1.39049\times 10^{-14}$ $9.38594\times 10^{-15}$
SVD $2.95307\times 10^{-15}$ $2.77759\times 10^{-15}$ $3.37643\times 10^{-15}$
LU $1.85695\times 10^{-14}$ $2.28425\times 10^{-14}$ $1.89852\times 10^{-14}$
@@
Table 1: Closeness of fit for each coordinate of the Lorenz system. The parameters are the same as those used for Fig.˜1.
Nevertheless, the situation drastically changes for $k=2$. For the same initial condition used before, this hyperparameter choice, common in NGRC experiments [10, 15, 12], drives the model into an ill-conditioned regime, with $\kappa(\Psi)=7.72\times 10^{16}$ (essentially the inverse of machine precision). In this setting, the numerical algorithms diverge as shown by the closeness of fit for each algorithm in Table˜2. While the SVD algorithm still produces solutions with high fitting accuracy, the other algorithms perform poorly. Specifically, the Cholesky algorithm fails, as the matrix becomes numerically singular. The LU algorithm produces invalid solutions because the $\arcsin$ function argument was larger than 1.
Algorithm $\theta_{x}$ $\theta_{y}$ $\theta_{z}$
Cholesky - - -
SVD $4.11460\times 10^{-15}$ $3.05268\times 10^{-15}$ $1.14009\times 10^{-14}$
LU - $0.70344$ -
Table 2: Closeness of fit for each coordinate of the Lorenz system. The symbol - indicates that the value could not be computed. The parameters are delay dimension $k=2$, and the rest of the parameters are the same as those used for Fig.˜1.
To reduce these issues, the regularizer parameter is varied as shown in Fig.˜3. For better visualization, we introduce the maximum closeness of fit over all Lorenz coordinates as $\theta_{\max}=\max\{\theta_{x},\theta_{y},\theta_{z}\}$. The top panel shows that increasing $\beta$ gradually reduces the condition number to $\mathcal{O}(10^{8})$, but also the fitting accuracy, inherent to the regularization trade-off. Small values of $\beta$ restore the existence of Cholesky solutions and bring their fitting accuracy closer to that of SVD solutions, although LU remains less accurate. By increasing $\beta$, all three algorithms have fitting performance that coincide at $\beta=10^{-10}$, and converge to similar solutions, as shown in the bottom panel in Fig.˜3.
Figure 3: Regularization reduces the difference among numerical algorithms during NGRC training. The top panel shows the maximum closeness of fit for Cholesky (dark gray), SVD (gray), and LU (light gray) solvers plotted together with the regularized condition number of $\Psi$ (red) with respect to the regularizer parameter $\beta$. All numerical algorithms attain similar fitting accuracy at $\beta=10^{-10}$. While the fitting performance gets worse as the regularizer parameter is increased, the bottom panel shows that the maximum pairwise difference among the solutions decays monotonically. The dots are the median, and bars are $75\%$ and $25\%$ quantiles over 50 different initial conditions. The parameters are $h=0.01$, delay dimension $k=2$, time lag $\tau=1$, the maximum degree $p=2$, and $N_{\mathrm{train}}=5000$.
This trade-off between fitting accuracy and numerical stability extends to the testing phase. We select a few regularizers used during training and contrast the results of Fig.˜3 with the NGRC’s resulting dynamics over multiple time steps, see Fig.˜4. The NGRC performance during testing is quantified by topological and statistical metrics, as described in detail in Section˜VII.1: valid prediction time (VPT) [39] quantifies for how long, in Lyapunov times, the NGRC trajectory is close to the original one; the distance between the induced successive local maxima map; and divergence error $E$ between the power spectrum densities estimated from the trajectories. Note that the columns of $\Psi$ are not normalized, as normalization can lead to different solutions [27]. Whenever the algorithm failed to find a solution, $\mathbf{w}_{i}(\beta)=0$. Moreover, if the trajectory of the NGRC during testing was unbounded, the metric value is set to 1-1.
Figure 4: Regularization also reduces the difference among numerical algorithms during NGRC testing. Comparison between three different numerical algorithms: Cholesky, SVD, and LU as the regularizer parameter $\beta$ is increased under the testing phase. From top to bottom, panels show box plots with respect to bounded models over 50 different initial conditions: valid prediction time (VPT), the distance between the induced successive local maxima maps, and the error $E$ between the power spectrum density, respectively. The colored hashed areas correspond to: absent/small (green), large (purple) regularization, and level of regularization in which all algorithms perform similarly within statistical confidence (orange). Blank spaces correspond to unbounded NGRC models over all initial conditions.. The parameters are $h=0.01$, delay dimension $k=2$, time lag $\tau=1$, the maximum degree $p=2$, and $N_{\mathrm{train}}=5000$.
Starting from the right in Fig.˜4, the purple region corresponds to the regime of large $\beta$. The VPT decreases monotonically, where all three algorithms obtain a solution whose performance drops because the regularization is too strong. The distance between successive maps and the divergence error remain statistically indistinguishable across initial conditions, until they deteriorate due to excessive regularization. The orange region corresponds to the regularizer value $\beta=10^{-10}$ that all algorithms perform equally over all metrics, and coincides with the value at which all algorithms have similar fitting accuracy shown in Fig.˜3. This regime corresponds to sufficient regularization for the best performance in all three metrics. As opposed to a smaller $\beta$, in the green region, where the three algorithms yield different NGRC performances. Cholesky and LU algorithms exhibit discontinuous behavior as $\beta$ changes, whereas SVD varies smoothly. SVD is robust against the ill-conditioning of $\Psi$. Surprisingly, at $\beta=0$, its NGRC performance remains unaltered. Thus, for this example, regularization is unnecessary for the SVD algorithm.
@@
Before proceeding, another benefit of studying the Lorenz example is that it sheds light on the relation between a large condition number and the autonomous NGRC dynamics. For a small time step $h$, when training NGRC, the task reduces to finding the governing equations of a dynamical system directly from time series data [40, 41, 42]. This gives an intuitive perspective into the autonomous NGRC dynamics during testing. Even small modeling errors — especially for terms that are absent in the true system — can cause instability when fitting an ODE from data, as already identified in [43]. For example, consider we attempt to learn $\dot{x}=-x,\dot{y}=0$ but instead recover $\dot{x}=-x,\dot{y}=\varepsilon y$. Although $\varepsilon$ may be small, the trajectory of the learned system diverges due to exponential growth in $y$. Similarly, if $\Psi$ is ill-conditioned, the estimated readout weights may differ substantially from the true ones, leading to unstable NGRC dynamics [26].
III Problem statement
The Lorenz example generated by the explicit Euler method demonstrates that the condition number $\kappa(\Psi)$ determines the accuracy of the NGRC model during training. In the ill-conditioned regime, the NGRC performance becomes highly sensitive to the numerical algorithm employed to solve the minimization problem in Eq.˜6. While regularization mitigates the effect of ill-conditioning, it introduces an extra hyperparameter, whose selection is non-trivial and has been studied extensively [36]. In particular, for forecasting tasks, the common approach is to train the model to predict one step forward, and then select the regularizer parameter that ensures accurate prediction over multiple time steps ahead. Consequently, finding the appropriate $\beta$ inevitably runs into extra computational cost — a cost that, for certain hyperparameter choices, could potentially be avoided. As shown using the SVD algorithm in the above example, NGRC can achieve accurate forecasts without any regularization.
This raises a fundamental question: can appropriate hyperparameter choices eliminate the need for regularization? To address this, we thoroughly investigate how the condition number varies with the hyperparameters: the delay dimension $k$, the time lag $\tau$, and the length of the training data $N_{\mathrm{train}}$. Our focus here is to investigate under which conditions $\Psi$ becomes ill-conditioned. We observe that $\Psi$ possesses rich internal structure: for $k>1$, it is simultaneously Vandermonde-like and Hankel-like because each entry corresponds to a monomial evaluated along a time-delayed trajectory of the original dynamics. As we will discuss, analyzing the conditioning of $\Psi$ is non-trivial and requires combining tools from numerical linear algebra with ergodic properties of dynamical systems. Then, we utilize this mapping as a guide for better choices of hyperparameters, enabling accurate NGRC models without the need for regularization.
IV The structure of $\Psi$ characterizes its conditioning